# Parametric Expensive Multi-Objective Optimization via Generative Solution Modeling

### v2

@@ 3.5 Alternating Optimization Loop
We maintain datasets per task-preference pair $(\boldsymbol{\lambda},\boldsymbol{\theta}_{k})$, storing top $Q\%$ solutions by scalarization score $\mathfrak{s}_{\boldsymbol{\lambda}}(F(\mathbf{x},\boldsymbol{\theta}_{k}))$ [tpe]. Preference vectors $\boldsymbol{\lambda}$ are generated via random sampling from $\Lambda$. Each elite solution $\mathbf{x}$ pairs with conditioning vector $\boldsymbol{c}=(\boldsymbol{\lambda},\boldsymbol{\theta}_{k})$ for generative model training.
@@
Method DTLZ-1 DTLZ-2 DTLZ-3 LAMP SOLAR MAGNETIC UAV
ST-MOBO 1.82e04 (4.80e03) 2.65e00 (6.30e-02) 4.74e02 (9.17e02) 5.80e-01 (1.20e-02) 6.81e-01 (2.50e-02) 8.70e-02 (2.20e-02) 4.31e-01 (2.00e-02)
qPAREGO 1.68e04 (4.03e03) 2.74e00 (5.30e-02) 4.28e02 (5.30e02) 5.79e-01 (1.30e-02) 6.64e-01 (2.20e-02) 1.06e-01 (2.10e-02) 4.27e-01 (1.80e-02)
qEHVI 3.23e04 (5.06e01) 2.33e00 (6.20e-02) 9.02e03 (1.90e-02) 6.14e-01 (3.00e-03) 6.07e-01 (1.30e-02) 2.18e-01 (1.20e-02) 4.89e-01 (3.00e-03)
PSL-MOBO 3.00e04 (3.52e03) 2.38e00 (4.90e-02) 9.22e03 (3.86e03) 5.42e-01 (1.40e-02) 5.96e-01 (4.00e-03) 7.40e-02 (1.80e-02) 4.03e-01 (1.50e-02)
PMT-MOBO 1.71e04 (4.70e03) 2.77e00 (5.50e-02) 9.44e02 (1.62e03) 5.83e-01 (1.20e-02) 7.10e-01 (1.90e-02) 1.73e-01 (3.00e-02) 4.78e-01 (1.30e-02)
PMT-MOBO-VAE 3.10e04 (2.22e03) 2.82e00 (1.48e-01) 5.29e03 (4.73e03) 6.02e-01 (2.50e-02) 7.17e-01 (3.80e-02) 1.49e-01 (5.10e-02) 4.93e-01 (3.10e-02)
PMT-MOBO-DDPM 3.25e04 (2.65e02) 2.70e00 (6.10e-02) 1.03e04 (1.84e03) 6.48e-01 (5.00e-03) 7.00e-01 (1.70e-02) 2.01e-01 (2.50e-02) 4.77e-01 (1.40e-02)
Table 1: Average Hypervolume Results on Sampled EMOP tasks in Synthetic and Real-World Benchmarks.
4 Results
We provide empirical experiments demonstrating our method’s ability to build effective inverse models for P-EMOP. Our experiments verify three key claims: (1) PMT-MOBO achieves faster convergence by exploiting inter-task synergies, (2) conditional generative models further improve solution quality, and (3) the built inverse model $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ can directly predict optimized solutions for unseen EMOP tasks without additional evaluations. We evaluate all methods using the hypervolume indicator [hv] across both seen and unseen EMOP tasks.
Benchmarks.
We evaluate our method on classical synthetic benchmarks: DTLZ-1, DTLZ-2, and DTLZ-3 [DTLZ], and real-world problems including lamp generation, solar rooftop generation [moo-gamut], magnetic sifter design [extremo], and UAV controller design [bicopter]. We modify these base EMOP problems so that each benchmark defines a family of related EMOPs, where $\boldsymbol{\theta}$ represents operational conditions (e.g., focal heights of lamps, building orientations of rooftops, and physical configurations of UAVs). More problem details are provided in the Appendix.
Baselines.
Our empirical studies include state-of-the-art MOBO methods: ST-MOBO [parego-2], qPAREGO [qpargeo], qEHVI [ehvi-3], and PSL-MOBO [moo-psl]. We also evaluate variants of our approaches that serve as ablation studies: PMT-MOBO (acquisition-driven only), PMT-MOBO-VAE, and PMT-MOBO-DDPM (full framework with different instantiated generative models).
Experimental Setup.
We optimize $K=8$ parameterized EMOPs simultaneously per benchmark with uniformly sampled task parameters. Each task uses 20 random initial solutions and the budget $T=50$, repeated over $U=20$ independent runs. We set $Q=10$ for elite solution extraction following [tpe] and $\beta$ follows [beta-schedule]. Additional details are in the Appendix.
Discussions.
Figure 2, Table 1 and Table 2 demonstrate our framework’s effectiveness on synthetic and real-world benchmarks. PMT-MOBO-VAE and PMT-MOBO-DDPM generally outperform state-of-the-art single-task MOBO baselines in discovering high-quality Pareto solutions across distinct tasks. It can be seen accordingly that both PMT-MOBO-X variants can outperform the compared baseline methods across nearly all the problem sets. The exception in magnetic sifter motivates us to further explore the integration of our methods with EHVI based acquisition functions so that the proposed framework can be more flexibly adapt to diverse applications. Table 2 further shows that our inverse models generalize well to unseen EMOP tasks, outperforming PSL-MOBO (the only baseline building inverse models post-optimization). To evaluate this generalization capability, we extract the trained inverse models after optimization and apply them to $W=100$ completely new tasks. For each unseen task, we query the inverse model with $S$ uniformly sampled preference vectors ($S=100$ for bi-objective, $S=1000$ for tri-objective) to generate predicted solutions, then compute the hypervolume of the resulting solution set.
Competitive Optimization Performance.
Figure 2 shows that PMT-MOBO-VAE and PMT-MOBO-DDPM generally outperform baselines in almost all benchmarks. PMT-MOBO-DDPM achieves best performance on DTLZ-1, DTLZ-3, and lamp generation, while PMT-MOBO-VAE leads on DTLZ-2, solar rooftop generation, and UAV controller design. The exception is low-dimensional magnetic sifter design, where qEHVI’s efficiency in identifying high-impact points under tight budgets proves advantageous, motivating future EHVI integration [ehvi-3, ehvi-4].
Effective Inverse Model Generalization.
Table 2 confirms our alternating optimization with generative modeling produces effective inverse models for unseen tasks. PMT-MOBO-VAE shows superior generalization on DTLZ-2 and solar rooftop generation, while PMT-MOBO-DDPM excels on DTLZ-1, DTLZ-3, lamp generation, and UAV controller design. We omit the magnetic sifter design problem set due to its computational costs of evaluating hundreds of tasks, where each task requires thousands of evaluations. Strong inverse model generalization typically correlates with effective multi-task optimization (e.g., PMT-MOBO-VAE on DTLZ-2, PMT-MOBO-DDPM on DTLZ-1/3). However, on UAV controller design, PMT-MOBO-VAE achieves the best optimization but PMT-MOBO-DDPM shows superior inverse model generalization, likely due to the intrinsic risk of posterior collapse issues when training VAE [lagging]. To further validate the quality of inverse models, Figure 3 visualizes the approximated Pareto fronts obtained by querying the inverse models with different task parameters and preference vectors. For the bi-objective solar rooftop generation problems, both PMT-MOBO variants discover superior Pareto fronts compared to PSL-MOBO. However, PMT-MOBO-DDPM produces less smooth Pareto fronts due to potential sampling instability inherent in diffusion models [reflected]. Similarly, for the tri-objective lamp generation problems, both PMT-MOBO variants significantly outperform the single-task PSL-MOBO baseline in terms of Pareto front quality and coverage. The specific task parameters used for each problem instance are indicated in Figure 3, demonstrating our method’s ability to adapt across diverse operational conditions.
Method DTLZ‑1 DTLZ‑2 DTLZ‑3 LAMP SOLAR UAV
PSL‑MOBO 2.22e4 (2.66e3) 1.90e0 (1.30e‑1) 1.07e4 (1.69e3) 4.38e‑1 (5.33e‑2) 3.67e‑1 (3.19e‑2) 4.12e‑1 (7.30e‑3)
PMT‑MOBO‑VAE 2.81e4 (3.52e3) 3.13e0 (1.18e‑2) 8.20e3 (4.26e3) 6.69e‑1 (7.20e‑2) 7.43e‑1 (1.65e‑2) 4.50e‑1 (8.00e‑3)
PMT‑MOBO‑DDPM 3.39e4 (7.22e2) 2.83e0 (7.10e‑2) 1.46e4 (2.41e3) 6.72e‑1 (6.06e‑2) 6.63e‑1 (2.80e‑2) 4.51e‑1 (2.20e‑2)
Table 2: Inverse Model Generalization Performance on Unseen EMOP Tasks in Synthetic and Real‑World Benchmarks.
Ablation Studies
Figure 2 demonstrates that PMT-MOBO generally outperforms ST-MOBO across nearly all benchmarks, empirically verifying Theorem 2 and confirming that leveraging inter-task synergies enables faster convergence. The superior performance of PMT-MOBO-VAE and PMT-MOBO-DDPM compared to PMT-MOBO highlights the value of integrating generative modeling. However, on diversity-oriented benchmarks like DTLZ-2, PMT-MOBO-DDPM may underperform due to training instability on broader variable distributions [DTLZ].
@@ 5.1 Inverse Modeling for Multi-Objective Optimization
Learning direct inverse maps from task–preference inputs to Pareto-optimal solutions enables flexible preference articulation for a family of expensive multi-objective optimization problems without expensive re-optimization. Early work treated the Pareto front as a smooth manifold to be approximated [Flemming]. Subsequently, [7018980] first integrated inverse modeling into the optimization loop itself, followed by Pareto Set Learning for EMOP [moo-psl], which employed inverse predictions for both post-hoc preference articulation and active search assistance. Recognizing the inherent knowledge in discovered Pareto-optimal solutions, subsequent work leveraged inverse models to streamline related multi-objective optimization (MOO) tasks: providing initial solutions for dynamic MOO [9440867], transferring quality solutions to related MOO tasks [inverse-2], or directly building inverse multi-task models [moo-finv] to learn multiple Pareto Sets or Pareto Fronts simultaneously.
@@ 5.2 Few-Shot Bayesian Optimization
Few-shot BO aims to rapidly adapt optimization configurations to new instances with minimal evaluations. Existing approaches focus on transferring optimizer components: multi-task GP kernels model inter-task relationships [swersky2013multi], meta-learned deep kernels configure GP surrogates from metadata of other optimization tasks [wistuba2021fewshotbayesianoptimizationdeep], reinforcement learning adapts acquisition functions [NEURIPS2021_3fab5890], and meta-priors enable transfer across heterogeneous spaces [JMLR:v25:23-0269, fan2024transfer].
@@ Appendix A Proofs of Theorem 2
Consider the conditional Gaussian distribution [GPML, TLBO], the conditional covariance matrix can be written as:
@@
where it can be verified that $\mathbf{\mathbf{K}}^{(m)}_{{\rm st},k^{*}}=\mathbf{K}^{(m)}_{k^{*},k^{*}}$ due to the assumption $\widetilde{\kappa}_{m}\bigl((\mathbf{x},\boldsymbol{\theta}_{k}),(\mathbf{x}^{\prime},\boldsymbol{\theta}_{k})\bigr)=\kappa_{m}(\mathbf{x},\mathbf{x}^{\prime}),\forall k\in[K]$. Since the kernel matrix, $\mathbf{K}^{m}_{\rm pmt}$, in equation (24), is positive semi-definite (PSD) and the matrix $\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I}$ in equation (28) is invertible, then according to the Schur complement theorem [schur], it holds in equation (28) that $\mathbf{K}^{(m)}_{k^{*},k^{*}}-B^{T}(\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I})^{-1}B$ and $\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I}$ are also both PSD. Considering the Minkowski determinant inequality that for positive semi-definite matrices $C$ and $D$, we have $|C+D|\geq|C|+|D|\geq|C|$. After substituting $C$ and $D$ by $C=\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I}$ and $D=\mathbf{K}^{(m)}_{k^{*},k^{*}}-B^{T}(\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I})^{-1}B$, respectively, the follows can be obtained:

### v3

@@ 3.5 Alternating Optimization Loop
We maintain datasets per task-preference pair $(\boldsymbol{\lambda},\boldsymbol{\theta}_{k})$, storing top $Q\%$ solutions by scalarization score $\mathfrak{s}_{\boldsymbol{\lambda}}(F(\mathbf{x},\boldsymbol{\theta}_{k}))$ [40]. Preference vectors $\boldsymbol{\lambda}$ are generated via random sampling from $\Lambda$. Each elite solution $\mathbf{x}$ pairs with conditioning vector $\boldsymbol{c}=(\boldsymbol{\lambda},\boldsymbol{\theta}_{k})$ for generative model training.
Method DTLZ-1 DTLZ-2 DTLZ-3 LAMP SOLAR MAGNETIC UAV
ST-MOBO 1.82e+04 (4.80e+03) 2.65e+00 (6.30e-02) 2.24e+03 (3.03e+03) 5.80e-01 (1.20e-02) 6.81e-01 (2.51e-02) 8.72e-02 (2.15e-02) 4.31e-01 (1.99e-02)
qPAREGO 1.68e+04 (4.03e+03) 2.74e+00 (5.30e-02) 1.42e+03 (1.48e+03) 5.79e-01 (1.30e-02) 6.64e-01 (2.21e-02) 1.06e-01 (2.05e-02) 4.27e-01 (1.82e-02)
qEHVI 3.23e+04 (5.06e+01)+- 2.33e+00 (6.20e-02) 2.66e+04 (1.85e-02)+- 6.14e-01 (3.00e-03) 6.07e-01 (1.29e-02) 2.18e-01 (1.15e-02)++ 4.89e-01 (2.60e-03)
PSL-MOBO 3.00e+04 (3.52e+03)≈- 2.38e+00 (4.90e-02) 2.31e+04 (6.44e+03)≈- 5.42e-01 (1.40e-02) 5.96e-01 (3.60e-03) 7.42e-02 (1.79e-02) 4.03e-01 (1.46e-02)
PPSL-MOBO 3.20e+04 (1.72e+03)+- 2.86e+00 (3.00e-01)≈- 2.71e+04 (7.89e+02)+- 6.09e-01 (1.61e-02) 6.35e-01 (1.15e-02) 9.41e-02 (1.69e-02) 5.85e-01 (3.19e-02)≈-
OCMOBO 1.28e+04 (3.04e+03) 2.99e+00 (2.12e-02)≈- 6.20e+02 (6.29e+02) 6.13e-01 (6.60e-03) 7.53e-01 (1.36e-02) 1.23e-01 (3.95e-02) 5.26e-01 (1.09e-02)
PMT-MOBO 1.71e+04 (4.70e+03) 2.77e+00 (5.50e-02) 3.78e+03 (5.40e+03) 5.83e-01 (1.22e-02) 7.10e-01 (1.90e-02) 1.73e-01 (2.99e-02)+- 4.78e-01 (1.27e-02)
PMT-MOBO-VAE 3.10e+04 (3.41e+03) 3.05e+00 (9.58e-02) 2.10e+04 (5.33e+03) 6.21e-01 (1.73e-02) 7.67e-01 (2.01e-02) 1.49e-01 (5.07e-02) 5.95e-01 (3.33e-02)
PMT-MOBO-DDPM 3.37e+04 (8.35e+02) 3.13e+00 (1.62e-02) 3.09e+04 (1.67e+03) 6.48e-01 (5.10e-03) 7.87e-01 (7.80e-03) 2.01e-01 (2.50e-02) 6.07e-01 (7.50e-03)
Table 1: Average Hypervolume results (mean $\pm$ std) across synthetic and real-world benchmarks. Statistical significance is assessed using the Wilcoxon signed-rank test at the 90% confidence level. Superscripts $^{s_{1}s_{2}}$ indicate that the competitor is significantly worse (-), significantly better (++), or statistically similar ($\approx$) relative to PMT-MOBO-VAE ($s_{1}$) and PMT-MOBO-DDPM ($s_{2}$).
4 Results
We provide empirical experiments demonstrating our method’s ability to build effective inverse models for P-EMOP. Our experiments verify three key claims: (1) PMT-MOBO achieves faster convergence by exploiting inter-task synergies, (2) conditional generative models further improve solution quality, and (3) the built inverse model $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ can directly predict optimized solutions for unseen EMOP tasks without additional evaluations. We evaluate all methods using the hypervolume indicator [67] across both seen and unseen EMOP tasks.
Benchmarks.
We evaluate our methods on classical synthetic benchmarks: DTLZ-1, DTLZ-2, and DTLZ-3 [13], and real-world problems including lamp generation, solar rooftop generation [35], magnetic sifter design [33], and UAV controller design [52]. We modify these base EMOP problems so that each benchmark defines a family of related EMOPs, where $\boldsymbol{\theta}$ represents operational conditions (e.g., focal heights of lamps, building orientations of rooftops, and physical configurations of UAVs). More problem details are provided in the Appendix B.
Baselines.
Our empirical studies include state-of-the-art single-task MOBO methods: ST-MOBO [41], qPAREGO [1], qEHVI [11], PSL-MOBO [32], and CDM-PSL [30], a diffusion model based MOBO variant. Moreover, to evaluate performance in utilizing inter-task synergies, we compare against multi-task MOBO methods: OCMOBO [6] and PPSL-MOBO [8], a parametric extension of PSL-MOBO [32]. Finally, we evaluate variants of our approaches that serve as ablation studies: PMT-MOBO (acquisition-driven only), PMT-MOBO-VAE, and PMT-MOBO-DDPM (full framework with different instantiated generative models).
Experimental Setup.
We optimize $K=8$ parameterized EMOPs simultaneously per benchmark with uniformly sampled task parameters. Each task uses 20 random initial solutions and the budget $T=50$, repeated over $U=20$ independent runs. We set $Q=10$ for elite solution extraction following [40] and $\beta$ follows [23]. Additional implementation details are in the Appendix C.
Discussions.
Figure 2, Table 1 and Table 2 demonstrate our framework’s effectiveness on synthetic and real-world benchmarks. PMT-MOBO-VAE and PMT-MOBO-DDPM generally show superior or comparable results compared to state-of-the-art single-task and multi-task MOBO baselines in discovering high-quality Pareto solutions across distinct tasks. The exception in magnetic sifter motivates us to further explore the integration of our methods with EHVI based acquisition functions so that the proposed framework can be more flexibly adapt to diverse applications. Table 2 further shows that our inverse models generalize well to unseen EMOP tasks, outperforming the single-task and multi-task inverse models constructed by PSL-MOBO and PPSL-MOBO. To evaluate this generalization capability, we extract the trained inverse models after optimization and apply them to $W=100$ completely new tasks. For each unseen task, we query the inverse model with $S$ uniformly sampled preference vectors ($S=100$ for bi-objective, $S=1000$ for tri-objective) to generate predicted solutions, then compute the hypervolume of the resulting solution set.
Competitive Optimization Performance.
Table 1 summarizes the hypervolume results across all benchmarks. Wilcoxon signed-rank tests (at 90% confidence) demonstrate that our proposed methods significantly outperform the baselines in most scenarios. Specifically, against the seven competing methods across all tasks (49 pairwise comparisons), PMT-MOBO-VAE attains a win/tie/loss record of 38/5/6, effectively dominating difficult tasks such as LAMP and SOLAR. The PMT-MOBO-DDPM further enhances optimization capability, achieving a record of 48/0/1, statistically surpassing all baselines in every benchmark except for the low-dimensional magnetic sifter design.
Effective Inverse Model Generalization.
Table 2 confirms that our alternating optimization strategy yields robust inverse models for unseen tasks. PMT-MOBO-DDPM demonstrates superior generalization across the majority of benchmarks, achieving the highest hypervolume on DTLZ-1, DTLZ-2, DTLZ-3, and LAMP generation. Notably, it statistically outperforms both PSL-MOBO and PPSL-MOBO baselines on every tested problem instance. PMT-MOBO-VAE excels on the remaining real-world engineering tasks, with the best performance on SOLAR and UAV controller design, though it struggles with the highly multimodal landscape of DTLZ-3 compared to the DDPM variant. We omit the magnetic sifter design set due to the prohibitive computational cost of evaluating hundreds of distinct tasks. To further validate these metrics, Figure 3 visualizes the approximated Pareto fronts. Both PMT-MOBO variants discover fronts with significantly better convergence and coverage than the PSL-MOBO baseline. Consistent with Table 2, PMT-MOBO-DDPM generally provides the broadest coverage, while PMT-MOBO-VAE offers smoother approximations, illustrating a potential trade-off between the expansive generation of diffusion models and the structured latent embedding of VAEs.
@@ 4 Results
Method DTLZ-1 DTLZ-2 DTLZ-3 LAMP SOLAR UAV
PSL-MOBO 2.82e+4 (1.78e+3) 1.90e+0 (1.30e-1) 2.39e+4 (3.22e+3)+- 5.01e-1 (7.32e-2) 3.54e-1 (4.06e-2) 4.12e-1 (7.30e-3)
PPSL-MOBO 2.24e+4 (1.10e+2) 2.98e+0 (3.56e-2) 2.67e+4 (6.08e+2)+- 6.12e-1 (1.05e-1)≈- 4.86e-1 (1.11e-1) 5.24e-1 (1.71e-1)
PMT-MOBO-VAE 2.88e+4 (2.57e+3) 3.10e+0 (2.99e-2) 1.78e+4 (7.28e+3) 6.31e-1 (7.33e-2) 7.93e-1 (1.56e-2) 6.08e-1 (1.88e-2)
PMT-MOBO-DDPM 3.47e+4 (4.85e+2) 3.13e+0 (1.41e-2) 3.20e+4 (1.98e+3) 6.70e-1 (5.82e-2) 7.65e-1 (1.30e-2) 6.05e-1 (2.44e-2)
Table 2: Inverse Model Generalization Performance on Unseen EMOP Tasks. Statistical significance is assessed using the Wilcoxon signed-rank test at the 90% confidence level. Superscripts $^{s_{1}s_{2}}$ indicate that the competitor is significantly worse (-), significantly better (++), or statistically similar ($\approx$) relative to PMT-MOBO-VAE ($s_{1}$) and PMT-MOBO-DDPM ($s_{2}$).
Ablation Studies and Sensitivity Analysis
To validate the necessity of our alternating optimization strategy, we compared the full framework against decoupled baselines (pure acquisition search and pure generative sampling). The results, detailed in Appendix D, confirm that neither component is sufficient in isolation. Specifically, generative sampling enhances the diversity of candidate generation, while the alternating acquisition search lowers the variance inherent in pure generative sampling, resulting in robust convergence. Moreover, figure 2 demonstrates the superior performance of PMT-MOBO-VAE and PMT-MOBO-DDPM compared to PMT-MOBO, also highlighting the value of integrating generative modeling.
Furthermore, our sensitivity analysis on the elite ratio parameter $Q$ reveals a critical trade-off: too large a $Q$ reduces selective pressure, pushing the generative model to learn low-quality solutions, whereas too small a $Q$ incurs large variance and provides insufficient data to train a well-behaved model. Empirically, $Q=10\%$ achieves the optimal balance across both VAE and DDPM variants (see Tables 6 and  7 in Appendix.D).
@@ 5.1 Inverse Modeling for Multi-Objective Optimization
Learning direct inverse maps from task–preference inputs to Pareto-optimal solutions enables flexible preference articulation for a family of expensive multi-objective optimization problems without expensive re-optimization. Early work treated the Pareto front as a smooth manifold to be approximated [17]. Subsequently, [9] first integrated inverse modeling into the optimization loop itself, followed by Pareto Set Learning for EMOP [32], which employed inverse predictions for both post-hoc preference articulation and active search assistance. Recognizing the inherent knowledge in discovered Pareto-optimal solutions, subsequent work leveraged inverse models to streamline related multi-objective optimization (MOO) tasks: providing initial solutions for dynamic MOO [64], transferring quality solutions to related MOO tasks [51], or directly building inverse multi-task models [57] to learn multiple Pareto Sets or Pareto Fronts simultaneously.
@@ 5.2 Few-Shot Bayesian Optimization
Few-shot BO aims to rapidly adapt optimization configurations to new instances with minimal evaluations. Existing approaches focus on transferring optimizer components: multi-task GP kernels model inter-task relationships [50], meta-learned deep kernels configure GP surrogates from metadata of other optimization tasks [60], reinforcement learning adapts acquisition functions [22], and meta-priors enable transfer across heterogeneous spaces [56, 15].
@@ Appendix A Proofs of Theorem 2
Consider the conditional Gaussian distribution [59, 46], the conditional covariance matrix can be written as:
@@
where it can be verified that $\mathbf{\mathbf{K}}^{(m)}_{{\rm st},k^{*}}=\mathbf{K}^{(m)}_{k^{*},k^{*}}$ due to the assumption $\widetilde{\kappa}_{m}\bigl((\mathbf{x},\boldsymbol{\theta}_{k}),(\mathbf{x}^{\prime},\boldsymbol{\theta}_{k})\bigr)=\kappa_{m}(\mathbf{x},\mathbf{x}^{\prime}),\forall k\in[K]$. Since the kernel matrix, $\mathbf{K}^{m}_{\rm pmt}$, in equation (24), is positive semi-definite (PSD) and the matrix $\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I}$ in equation (28) is invertible, then according to the Schur complement theorem [63], it holds in equation (28) that $\mathbf{K}^{(m)}_{k^{*},k^{*}}-B^{T}(\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I})^{-1}B$ and $\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I}$ are also both PSD. Considering the Minkowski determinant inequality that for positive semi-definite matrices $C$ and $D$, we have $|C+D|\geq|C|+|D|\geq|C|$. After substituting $C$ and $D$ by $C=\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I}$ and $D=\mathbf{K}^{(m)}_{k^{*},k^{*}}-B^{T}(\mathbf{K}^{(m)}_{\setminus k^{*}}+\sigma_{\epsilon}^{-2}\mathbf{I})^{-1}B$, respectively, the follows can be obtained: